### v1

# Clairaut anti-invariant Riemannian map from/to Kähler manifolds admitting Ricci solitons

Abstract
The aim of this article is to study the Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons. We find the curvature relations and calculate the Ricci tensor under different conditions. We obtain conditions for the range and kernel of these maps to be Einstein. Next, we find the scalar curvature for range. Finally, we give some non-trivial examples of Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons.
M. S. C. 2020: 53B20, 53B35.
Keywords: Kähler manifolds, Riemannian maps, Clairaut Riemannian maps, anti-invariant Riemannian maps, Ricci solitons.
1 Introduction
The submersions and immersions play an important role in the study of smooth maps between smooth manifolds. In the theory of Riemannian manifolds, Riemannian maps were introduced by Fisher [4] in 1992 as a generalization of isometric immersion and Riemannian submersion. The notable property of Riemannian maps is that these maps satisfy the generalized eikonal equation, which connects the geometric and physical optics.
The geometry of Riemannian submersions was discussed by Falcitelli et al. [3]. As a generalization of holomorphic Riemannian submersions, Şahin [15] introduced holomorphic Riemannian maps. Further, as a generalization of anti-invariant Riemannian submersions, Şahin [14] introduced anti-invariant Riemannian maps from almost Hermitian manifolds to Riemannian manifolds.
The idea of Clairaut theorem is based on geodesics on the surface of revolution. This theorem states that for any geodesic $\gamma$($\gamma:I\subset R\rightarrow S$) on a surface of revolution, $rsin\theta$ is constant along $\gamma$, where $\theta(s)$ is the angle between $\gamma(s)$ and the meridian curve through $\gamma(s)$, $s\in I$ and $r:S\rightarrow R^{+}$ is the distance of a point of $S$ from the axis of rotation. Bishop [1] generalized this idea on submersion theory and defined Clairaut submersion. In [18], Şahin gave the idea of Clairaut Riemannian maps, in which he obtained the necessary and sufficient conditions for a Riemannian map to be a Clairaut Riemannian map. Recently, Meena et al. [7, 22] have studied Clairaut anti-invariant Riemannian maps from/to Kähler manifolds.
Ricci solitons are the well-known object in Riemannian geometry. It comes out as the solution of Einstein’s field equations within the framework of general relativity. Since the equation of Ricci soliton has relations with string theory, it has also become an attractive subject of study for physicists. [6] Ricci soliton appears during the analysis of the Ricci flow
$\frac{\partial g(t)}{\partial t}=-2g(t),$
where $g(t)$ is a one-parameter family of metrics on a manifold.
A smooth vector field $Z$ on a Riemannian manifold is called a conformal vector field if the Lie derivative of the metric g with respect to X can be expressed as a multiple of a function to the metric g, and this function is called the conformal factor of the conformal vector field $Z$ [2].
A connected Riemannian manifold $(M,g_{M})$ is called $\alpha$ Ricci soliton [5], denoted by $(M,g_{M},\xi,\lambda)$, if it satisfies the following equation
$\frac{1}{2}(L_{\xi}g)(X,Y)+\alpha Ric(X,Y)+\lambda g_{M}(X,Y)=0,$ (1.1)
where $\alpha,\lambda\in\operatorname{\mathbb{R}},$ $\xi$ is a potential vector field, $L_{\xi}g$ denotes the Lie derivative of metric tensor $g$ with respect to $\xi$, Ric denotes the Ricci curvature.
The $\alpha$ Ricci soliton is called shrinking, steady or expanding according as $\lambda<0,\lambda=0$ or $\lambda>0,$ respectively. If $\alpha=1,$ then (1.1) reduces to Ricci soliton. If $\xi$ is gradient of some function $f$ which is defined on $M$, i.e., $\xi=gradf$, $f\in C^{\infty}(M)$, then from (1.1), we get
$Hess^{f}(X,Y)+Ric(X,Y)+\lambda g(X,Y)=0,$
where $Hess^{f}$ is hessian form. Then $(M,g_{M},\nabla f,\lambda)$ is called gradient Ricci soliton and the function f is called potential function.
Pigola et al. [11] gave the idea of almost Ricci soliton by using $\lambda$ as a variable function rather than constant function. In this case $(M,g_{M},\xi,\lambda)$ is called an almost Ricci soliton.
Meriç and Kiliç [8] study about Riemannian submersions whose total manifolds admit a Ricci soliton. Yadav and meena [21] have studied Riemannian maps whose total admit a Ricci soliton. In [7, 23], Meena and yadav have studied Clairaut Riemannian maps whose total and base manifolds admit a Ricci soliton. Watson [20] studies Hermitian submersion whose total manifold admits Ricci soliton.
In this article, we study Clairaut anti-invariant Riemannian maps from/to Kähler manifolds admitting Ricci solitons. The article is divided into four sections. In section 2, we recall some definitions and results which are useful to current study. In section 3, we study Clairaut anti-invariant Rieannian maps from a Kähler manifold admitting Ricci soliton to a Riemannian manifold and give non trivial examples for the existence. In section 4, we study Clairaut anti-invariant Riemannian maps from a Riemannian manifold to a Kähler manifold admitting Ricci soliton and construct some non-trivial examples.
2 Preliminaries
Let $F$ be a smooth map between two Riemannian manifolds $(M,g_{M})$ and $(N,g_{N})$ of dimension $m,n,$ respectively. Let $\mathcal{V}_{p}=kerF_{*p}$ at $p\in M,$ denotes vertical distribution or kernel space of $F_{*}$ and $\mathcal{H}_{p}=(kerF_{*p})^{\perp}$ in $T_{p}M$ is the orthogonal complementary space of $\mathcal{V}_{p}$. Then the tangent space $T_{p}M$ at $p\in M$ has the decomposition,
$T_{p}M=(kerF_{*p})\oplus(kerF_{*p})^{\perp}=\mathcal{V}_{p}\oplus\mathcal{H}_{p}.$
Let the range of $F_{*}$ be denoted by $rangeF_{*p}$ at $p\in M$, and $(rangeF_{*p})^{\perp}$ be the orthogonal complementary space of $rangeF_{*p}$ in the tangent space $T_{F(p)}N$ of $N$ at $F(p)\in N$. Since $rankF\leq\min\{m,n\}$, therefore we get, $(rangeF_{*p})^{\perp}\neq\{0\}$. Thus, the tangent space $T_{F(p)}N$ of $N$ at $F(p)\in N$ has the following decomposition:
$T_{F(p)}N=(rangeF_{*p})\oplus(rangeF_{*p})^{\perp}.$
A smooth map $F$ is said to be a Riemannian map at $p\in M$ if the horizontal restriction $F^{h}_{p}:(kerF_{*p})^{\perp}\rightarrow(rangeF_{*p})$ is a linear isometry between the inner product spaces $((kerF_{*p})^{\perp},g_{N}{(p)}|(kerF_{*p})^{\perp})$ and $((rangeF_{*p},g_{N}(q)|(rangeF_{*p})),q=F(p)$. In other words, $F_{*}$ satisfies the equation [17]
${}g_{N}(F_{*}X,F_{*}Y)=g_{M}(X,Y)~{}~{}\forall X,Y\in\Gamma(kerF_{*})^{\perp}.$ (2.1)
For Riemannian submersion, O’Neill [10] defined fundamental tensor fields $\mathcal{T}$ and $\mathcal{A}$ by
$A_{E}{F}=\mathcal{H}\nabla_{HE}VF+\mathcal{V}\nabla_{HE}HF,\$ (2.2)
$T_{E}{F}=\mathcal{H}\nabla_{VE}VF+\mathcal{V}\nabla_{VE}HF,\$ (2.3)
where $E,F\in\Gamma(M),\nabla$ is the Levi-Civita connection of $g_{M}$ and $\mathcal{V},\mathcal{H}$ denote the projections to vertical subbundle and horizontal subbundle, respectively. For any $E\in\Gamma(TM)$, $T_{E}$ and $A_{E}$ are skewsymmetric operators on $(\Gamma(TM),g)$ reversing the horizontal and the vertical distributions. One can check easily that $\mathcal{T}$ is vertical, i.e., $\mathcal{T}_{E}=T_{\mathcal{V}E}$, and $\mathcal{A}$ is horizontal, i.e., $\mathcal{A}=A_{HE}$.
We observe that the tensor field $T$ is symmetric for vertical vector fields, i.e.,
$T_{U}W=T_{W}U~{}~{}\forall~{}~{}U,W\in\Gamma(kerF_{*}),\$
and tensor field $A$ is anti-symetric for horizontal vector fields i.e.,
$A_{X}Y=-A_{Y}X=\frac{1}{2}\mathcal{V}[X,Y]~{}~{}\forall~{}~{}X,Y\in\Gamma(kerF_{*})^{\perp}.\$
Using equation (2.2) and (2.3), we have the following lemma.
Lemma 2.1.
[10] Let $X,Y\in\Gamma(kerF_{*})^{\perp}$ and $V,W\in\Gamma(kerF_{*}).$ Then
$\nabla_{V}W=T_{V}W+\hat{\nabla}_{V}W,$
$\nabla_{V}X=\mathcal{H}\nabla_{V}X+T_{V}X,$
$\nabla_{X}V=\mathcal{A}_{X}V+\mathcal{V}\nabla_{X}V,$
$\nabla_{X}Y=\mathcal{H}\nabla_{X}Y+\mathcal{A}_{X}Y.$
If $X$ is a basic vector field. then $\mathcal{H}\nabla_{V}X=A_{X}V.$
For any vector field $X$ on $M$ and any section $V$ of $(rangeF_{*})^{\perp}$, we define the connection $\nabla^{F\perp}$ as the orthogonal projection of $\nabla_{X}V$ on $(rangeF_{*})^{\perp}.$ One can easily see that $\nabla^{F\perp}$ is a linear connection on $(rangeF_{*})^{\perp}$ with $\nabla^{F\perp}g_{N}=0.$ For a Riemannian map $F$, we have [16]
$\nabla^{N}_{F_{*}X}V=-S_{V}{F_{*}X}+\nabla^{F\perp}_{X}V,$ (2.4)
where $S_{V}{F_{*}X}$ is the tangential component and $\nabla^{F\perp}_{X}V$ is the orthogonal component of $\nabla^{N}_{F_{*}X}V$. It can be easily seen that $\nabla^{N}_{F_{*}X}V$ is obtained from the pullback connection of $\nabla^{N}$. Thus, at $p\in M$, we have $\nabla^{N}_{F_{*}X}V(p)\in T_{F(p)}N,S_{V}{F_{*}X}\in F_{*p}(T_{p}M)$ and $\nabla^{F\perp}_{X}V\in(F_{*p}(T_{p}M))^{\perp}$. It follows that $S_{V}{F_{*}X}$ is bilinear in $V$, $F_{*}X$ and $S_{V}{F_{*}X}$ at $p$ depends only on $V_{p}$ and $F_{*p}X_{p}$.
By direct computations, we obtain
$g_{N}(S_{V}{F_{*}X,{F_{*}Y}})=g_{N}\big{(}V,(\nabla F_{*})(X,Y)\big{)}\$
for all X, Y $\in\Gamma(kerF_{*})^{\perp}~{}and~{}V\in\Gamma(rangeF_{*})^{\perp}.$ Since $\nabla F_{*}$ is symmetric, it follows that $S_{V}$ is a symmetric linear transformation of $rangeF_{*}$.
 Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a smooth map between smooth manifolds. The second fundamental form of $F$ is the map [9]
$\nabla F_{*}:\Gamma(TM)\times\Gamma(TM)\rightarrow\Gamma_{F}(TN)\$
defined by
$(\nabla F_{*})(X,Y)=\nabla^{N_{F}}_{X}{F_{*}}Y-F_{*}(\nabla^{M}_{X}Y),$
where $\nabla^{M}$ is a linear connection on $M$.
Theorem 2.1.
[7] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Riemannian map between Riemannian manifolds such that $(rangeF_{*})^{\perp}$ is totally geodesic and $rangeF_{*}$ is connected, let $\beta$ and $\eta=Fo\beta$ be geodesics on $M$ and $N,$ respectively.Then $F$ is a Clairaut Riemannian map with $\tilde{s}=e^{g}$ if and only if any one of the following conditions holds:
  • (i)
    $S_{V}F_{*}X=-V(g)F_{*}X$, where $F_{*}X\in\Gamma(rangeF_{*})$ and $V\in\Gamma(rangeF_{*})^{\perp}$ are components of $\dot{\eta}(t)$.
  • (ii)
    $F$ is an umbilical map, and has $H=-\nabla^{N}g$ , where $g$ is a smooth function on $N$ and $H$ is the mean curvature vector field of $rangeF_{*}$.
Definition 2.1.
[7] A Riemannian map $F:(M,g_{M})\rightarrow(N,g_{N})$ between Riemannian manifolds is called a Clairaut Riemannian map if there is a function $s:N\rightarrow R^{+}$ such that for every geodesic $\eta$ on $N,$ the function $(so\eta)sin\omega(t)$ is constant, where $F_{*}X\in\Gamma(rangeF_{*})$ and $U\in\Gamma(rangeF_{*})^{\perp}$ are the vertical and horizontal components of $\dot{\eta}(t)$, and $\omega(t)$ is the angle between $\dot{\eta}(t)$ and $U$ for all t.
Lemma 2.2.
[13] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Riemannian map between Riemannian manifolds. Then $F$ is an umbilical Riemannian map if and only if
$(\nabla F_{*})(X,Y)=g_{M}(X,Y)H$
for all $X,Y\in\Gamma(kerF_{*})^{\perp}$ and $H$ is nowhere zero mean curvature vector field on $(rangeF_{*})^{\perp}$.
In [25], it is proved that an almost Hermitian manifold $(M,g_{M})$ admits a tensor $J$ of type (1,1) such that $J^{2}=-I$ and
$g_{M}(JX,JY)=g_{M}(X,Y)$ (2.5)
for all $X,Y\in\Gamma(TM).$ An almost Hermitian manifold $M$ is called a Kähler manifold if
$(\nabla_{X}J)Y=0\$ (2.6)
for all $X,Y\in\Gamma(TM)$, where $\nabla$ is Levi-Civita connection on $M$.
Theorem 2.2.
[25] For a Kähler manifold $M$, follwing properties are satisfied:
  • (i)
    R(X, Y) JZ = JR(X, Y)Z,
  • (ii)
    R(JX, JY) = R(X, Y),
  • (iii)
    Ric(X, Y) = $\frac{1}{2}\{Tr.JoR(X,JY)\},QJ=JQ,$
  • (iv)
    Ric(JX, JY) = Ric(X, Y),
where $X,Y,Z\in\Gamma(TM)$, $Q$ is the Ricci operator and $R$ is Riemannian curvature tensor.
Definition 2.2.
[14] Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Riemannian map from an almost Hermitian manifold to a Riemannian manifold. We say that $F$ is an anti-invariant Riemannian map if
$J(kerF_{*})\subset(kerF_{*})^{\perp}.$
We denote the orthogonal complementary subbundle to $J(kerF_{*})$ in $(kerF_{*})^{\perp}$ by $\mu$.
Next, let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be an anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold. Then, for any $X\in\Gamma((kerF_{*})^{\perp}),$ we have
$JX=BX+CX,$ (2.7)
where $BX\in\Gamma(kerF_{*})$ and $CX\in\Gamma(\mu).$
Theorem 2.3.
[18] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Riemannian map between Riemannian manifolds with connected fibers, then $F$ is a Clairaut Riemannian map with $\tilde{r}=e^{f}$ if and only if each fiber is umbilical and has mean curvature vector field $H=-gradf$, where $gradf$ is the gradient of the function $f$ with respect to $g_{M}$.
Definition 2.3.
[12] Let $F:(N,g_{N})\rightarrow(M,g_{M})$ be a proper Riemannian map from a Riemannian manifold to an almost Hermitian manifold with almost complex structure $J.$ We say that $F$ is an anti-invariant Riemannian map at $q\in N$ if $J(rangeF_{*}q)\subset(rangeF_{*q})^{\perp}.$
If $F$ is an anti-invariant Riemannian map for every $q\in N,$ then $F$ is called an anti-invariant Riemannian map. In this case, we denote the orthogonal subbundle to $J(rangeF_{*})$ in $(rangeF_{*})^{\perp}$ by $\nu$, i.e., $(rangeF_{*})^{\perp}=J(rangeF_{*})\oplus\nu.$
For any $E\in\Gamma(rangeF_{*})^{\perp},$ we have [19]
$JE=PE+QE,$ (2.8)
where $PE\in\Gamma(rangeF_{*})$ and $QE\in\nu$. Note that if $\nu=0$ then $F$ is called a Lagrangian Riemannian map.
Theorem 2.4.
[7] Let $F:(M,g_{M})\rightarrow(N,g_{N})$ be a Clairaut Riemannian map with $\tilde{s}=e^{g}$ between Riemannian manifolds such that $kerF_{*}$ is minimal. Then $F$ is harmonic if and only if g is a constant function on N.
3 Clairaut anti-invariant Riemannian maps from Kähler manifolds admitting Ricci soliton to Riemannian manifolds
In this section, we study Clairaut anti-invariant Riemannian maps from Kähler manifolds admitting Ricci solitons to Riemannian manifolds. Throughout this section, we are taking $dim(kerF_{*})>1.$ Below we provide a revised version of proposition (3.9) [22].
Proposition 3.1.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold with $\tilde{r}=e^{f}.$ Then, Ricci tensors of $(M,g_{M},J)$ are given by
$\begin{split}Ric(U,V)=Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+rg_{M}(\nabla_{JU}\nabla f,JV)-div\Big{(}A(JU,JV)\Big{)}\end{split}$ (3.1)
$\begin{split}Ric(U,X)=&-(r+1)g_{M}(\nabla_{BX}\nabla f,JU)+div\Big{(}A(JU,CX)\Big{)},\\ &-rg_{M}(\nabla_{JU}\nabla f,CX)+Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)\\ &+\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(X_{i},JU),BX)\end{split}$ (3.2)
and
$\begin{split}Ric(X,Y)&=Ric^{(KerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+\sum_{i=1}^{r+s}g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-rCX(f)CY(f)+\sum_{j=1}^{r}g_{M}(A_{CX}u_{j},A_{CY}u_{j})\\ &+div\Big{(}A(CX,CY)\Big{)}+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},CY),(\nabla F_{*})(CX,X_{i})\Big{)}+g_{N}\Big{(}(\nabla F_{*})(CX,CY),\tau^{(kerF_{*})^{\perp}}\Big{)}\\ &-(r+1)g_{M}(\nabla_{BX}\nabla f,CY)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(CY,X_{i}),BX\Big{)}\\ &-(r+1)g_{M}(\nabla_{BY}\nabla f,CX)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(CX,X_{i}),BY\Big{)}\end{split}$ (3.3)
for all $X,Y\in\Gamma(kerF_{*})^{\perp},U,V\in\Gamma(kerF_{*}),$ where $\{u_{1},u_{2},...,u_{r}\},\{X_{1},X_{2},...,X_{r+s}\}$ and $\{\mu_{1},\mu_{2},...,\mu_{s}\}$ are orthonormal frames of $kerF_{*},J(kerF_{*})\oplus\mu$ and $\mu,$ respectively, and Ric, $Ric^{(kerF_{*})}$ and $Ric^{rangeF_{*}}$ are Ricci tensors on $M$, $kerF_{*}$ and $rangeF_{*}$, respectively.
Proof.
Let $F$ be an anti-invariant Riemannian map from a Kähler manifold $(M,g_{M},J)$ to a Riemannian manifold $(N,g_{N}).$ In [14], Şahin defined $Ric(U,V)$ as
$\begin{split}Ric(U,V)=&\sum_{j=1}^{r}\Big{\{}-g_{M}\Big{(}\nabla_{JU}T)(u_{j},u_{j}),JV\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,JV),u_{j}\Big{)}\\ &+g_{M}(T_{u_{j}}JU,T_{u_{j}}JV)-g_{M}(A_{JU}u_{j},A_{JV}u_{j})\Big{\}}\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+g_{N}\Big{(}(\nabla F_{*})(JU,JV),\tau^{(kerF_{*})^{\perp}}\Big{)}\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},JV),(\nabla F_{*})(JU,X_{i})\Big{)}.\end{split}$ (3.4)
Since $F$ is a Clairaut anti-invariant Riemannian map, using clairaut condition $T_{U}V=-g(U,V)\nabla f,$ Lemma 3.7 [22] and Lemma 3.8 [22], we get
$-\sum_{j=1}^{r}g_{M}\Big{(}\nabla_{JU}T)(u_{j},u_{j}),JV\Big{)}=rg_{M}(\nabla_{JU}\nabla f,JV),$
$\sum_{j=1}^{r}g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,JV),u_{j}\Big{)}=div\Big{(}A(JU,JV)\Big{)},$
$\sum_{j=1}^{r}g_{M}(T_{u_{j}}JU,T_{u_{j}}JV)=0,$
$\sum_{j=1}^{r}g_{M}(A_{JU}u_{j},A_{JV}u_{j})=0,$
$g_{N}\Big{(}\nabla F_{*})(JU,JV),\tau^{(kerF_{*})^{\perp}}\Big{)}=0,$
$\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},JV),(\nabla F_{*})(JU,X_{i})\Big{)}=0.$
Therefore using above results in (3.4), we get
$Ric(U,V)=Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+rg_{M}(\nabla_{JU}\nabla f,JV)-div\Big{(}A(JU,JV)\Big{)},$
Next, in [14], $Ric(U,X)$ is defined as
$\begin{split}Ric(U,X)=&\sum_{j=1}^{r}\Big{\{}g_{M}\Big{(}(\nabla_{BX}T)(u_{j},u_{j}),JU\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}T)(BX,u_{j}),JU\Big{)}\\ &+g_{M}\Big{(}(\nabla_{JU}T)(u_{j},u_{j}),CX\Big{)}+g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,CX),u_{j}\Big{)}\\ &-g_{M}(T_{u_{j}}JU,T_{u_{j}}CX)+g_{M}(A_{JU}u_{j},A_{CX}u_{j})\Big{\}}+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}CX)\\ &+g_{N}\Big{(}(\nabla F_{*})(JU,CX),\tau^{(kerF_{*})^{\perp}}\Big{)}+\sum_{i=1}^{r+s}\Big{\{}g_{M}\Big{(}(\nabla_{X_{i}}A)(X_{i},JU),BX\Big{)}\\ &+2g_{M}(A_{X_{i}}JU,T_{BX}X_{i})-g_{N}\Big{(}(\nabla F_{*})(X_{i},CX),(\nabla F_{*})(JU,X_{i}))\Big{\}}.\end{split}$ (3.5)
Since $F$ is a Clairaut anti-invariant Riemanian map, using clairaut condition $T_{U}V=-g(U,V)\nabla f,$ lemma 3.7 [22] and lemma 3.8 [22], we get
$\begin{split}&\sum_{j=1}^{r}\Big{\{}g_{M}\Big{(}(\nabla_{BX}T)(u_{j},u_{j}),JU\Big{)}-g_{M}\Big{(}(\nabla_{u_{j}}T)(BX,u_{j}),JU\Big{)}+2g_{M}(A_{X_{i}}JU,T_{BX}X_{i})\Big{\}}\\ &=-(r+1)g_{M}(\nabla_{BX}\nabla f,JU),\end{split}$
$\sum_{j=1}^{r}g_{M}\Big{(}(\nabla_{u_{j}}A)(JU,CX),u_{j}\Big{)}=div\Big{(}A(JU,CX)\Big{)},$
$\sum_{j=1}^{r}g_{M}(T_{u_{j}}JU,T_{u_{j}}CX)=0,$
$\sum_{j=1}^{r}g_{M}(A_{JU}u_{j},A_{CX}u_{j})=0,$
$g_{N}\Big{(}(\nabla F_{*})(JU,X_{i}),(\nabla F_{*})(X_{i},CX)=0,$
$g_{N}\Big{(}(\nabla F_{*})(JU,CX),\tau^{(kerF_{*})^{\perp}}\Big{)}=0,$
$g_{M}\Big{(}(\nabla_{JU}T)(u_{j},u_{j}),CX\Big{)}=-rg_{M}(\nabla_{JU}\nabla f,CX).$
Using above results in (3.5), we get
$\begin{split}Ric(U,X)=&-(r+1)g_{M}(\nabla_{BX}\nabla f,JU)+div\Big{(}A(JU,CX)\Big{)}\\ &-rg_{M}(\nabla_{JU}\nabla f,CX)+Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)\\ &+\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(X_{i},JU),BX).\end{split}$
Next, Since $M$ is a Kähler manifold, we have
$Ric(X,Y)=Ric(JX,JY)~{}~{}\forall~{}~{}X,Y\in\Gamma(TM).$ (3.6)
Using (2.7), we have
$Ric(JX,JY)=Ric(BX,BY)+Ric(CX,CY)+Ric(BX,CY)+Ric(CX,BY).$ (3.7)
Next, since
$Ric(BX,BY)=\sum_{j=1}^{r}g_{M}(R(u_{j},BX)BY,u_{j})+\sum_{i=1}^{r+s}g_{M}(R(X_{i},BX)BY,X_{i}),$
using lemma 52 ([16]), we get
$\begin{split}Ric(BX,BY)&=Ric^{(BX,BY)}-rg_{M}(H,T_{BX}BY)+g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)\\ &+\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}T)_{BX}BY,X_{i}\Big{)}.\end{split}$
Since $F$ is a Clairaut anti-invariant Riemannian map, applying the Clairaut condition $T_{U}V=-g_{M}(U,V)\nabla f$, and metric compatibility condition of the Levi-civita connection $\nabla$, we get
$\begin{split}Ric(BX,BY)&=Ric^{(BX,BY)}-r|\nabla f|^{2}g_{M}(BX,BY)+\sum_{i=1}^{r+s}g_{M}(A_{X_{i}}BX,A_{X_{i}}BY)\\ &-g_{M}(BX,BY)div(\nabla f).\end{split}$ (3.8)
Similar computations provide the following three equations:
$\begin{split}Ric(CX,CY)=&-rg_{M}(\nabla_{CX}\nabla f,CY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-rCX(f)CY(f)+\sum_{j=1}^{r}g_{M}(A_{CX}u_{j},A_{CY}u_{j})+div\Big{(}A(CX,CY)\Big{)}\\ &-\sum_{i=1}^{r+s}g_{N}\Big{(}(\nabla F_{*})(X_{i},CY),(\nabla F_{*})(CX,X_{i})\Big{)}\\ &+g_{N}\Big{(}(\nabla F_{*})(CX,CY),\tau^{(kerF_{*})^{\perp}}\Big{)},\end{split}$ (3.9)
$\begin{split}Ric(BX,CY)=-(r+1)g_{M}(\nabla_{BX}\nabla f,CY)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)_{CY}{X_{i}},BX),\end{split}$ (3.10)
$\begin{split}Ric(CX,BY)&=-(r+1)g_{M}(\nabla_{BY}\nabla f,CX)-\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)_{CX}{X_{i}},BY).\end{split}$ (3.11)
Adding (3.8), (3.9), (3.10), (3.11), we obtain (3.7) and with the help of (3.6), we get (3.3). ∎
Corollary 3.1.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langragian Riemannian map from a Kähler manifold to a Riemannian manifold with $\tilde{r}=e^{f}.$ Then, Ricci tensors on $(M,g_{M})$ are given by
$\begin{split}Ric(U,V)=rg_{M}(\nabla_{JU}\nabla f,JV)+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV),\end{split}$ (3.12)
$\begin{split}Ric(U,X)=0,\end{split}$ (3.13)
and
$\begin{split}Ric(X,Y)=Ric^{(KerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY).\end{split}$ (3.14)
Proof.
In case of Langragian Riemannian map, $\mu=0$ or $CX=0=CY.$
Therefore (3.1), (3.2) and (3.3) reduce to (3.12), (3.13) and (3.14), respectively. This completes the proof. ∎
Corollary 3.2.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold to a Riemannian manifold with with $\tilde{r}=e^{f}$ such that $(kerF_{*})^{\perp}$ integrable. Then
$\begin{split}Ric(X,Y)&=Ric^{(kerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &-rg_{M}(\nabla_{CX}\nabla f,CY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)-rCX(f)CY(f),\end{split}$ (3.15)
where $X,Y\in\Gamma{(kerF_{*})^{\perp}}.$
Proof.
Since $F$ is a totally geodesic Clairaut anti-invariant Riemannian map, the second fundamental form of the map $\nabla F_{*}$ vanishes. Since $(kerF_{*})^{\perp}$ is integrable, we have $A_{X}=0~{}~{}\forall X\in\Gamma(kerF_{*})^{\perp}$. Therefore the equation (3.3), reduces to (3.15). ∎
Theorem 3.1.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold admitting Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(KerF_{*})^{\perp}$ is totally geodesic. Then $F_{*}(JkerF_{*})$ is an $\alpha$- Ricci soliton.
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$
$\frac{1}{2}(L_{\eta}g_{M})(U,V)+Ric(U,V)+\lambda g_{M}(U,V)=0~{}~{}\forall~{}~{}U,V\in\Gamma(KerF_{*}),$
which can be rewritten as
$\frac{1}{2}\Big{(}g_{M}(\nabla_{U}\eta,V)+g_{M}(\nabla_{V}\eta,U)\Big{)}+Ric(U,V)+\lambda g_{M}(U,V)=0.$
Making use of (3.1) in the above equation, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{U}\eta,V)+g_{M}(\nabla_{V}\eta,U)\Big{)}+rg_{M}(\nabla_{JU}\nabla f,JV)\\ &-div\Big{(}A(JU,JV)\Big{)}+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)\\ &+\lambda g_{M}(U,V)=0.\end{split}$ (3.16)
Since $(kerF_{*})^{\perp}$ is totally geodesic, $A_{X}=0~{}~{}\forall X\in\Gamma(kerF_{*})^{\perp}.$ Thus (3.16) reduces to
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{U}\eta,V)+g_{M}(\nabla_{V}\eta,U)\Big{)}+rg_{M}(\nabla_{JU}\nabla f,JV)+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)\\ &+\lambda g_{M}(U,V)=0.\end{split}$ (3.17)
Using (2.5) and (2.6) in (3.17), we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{U}J\eta,JV)+g_{M}(\nabla_{V}J\eta,JU)\Big{)}+rg_{M}(\nabla_{JU}\nabla f,JV)\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\lambda g_{M}(JU,JV)=0.\end{split}$ (3.18)
If $\eta=W\in\Gamma(kerF_{*})$ or in $\Gamma(\mu)$, then above equation gives
$\begin{split}&\frac{1}{2}\Big{(}g(A_{J\eta}U,JV)+g(JU,A_{J\eta}V)\Big{)}+rg_{M}(\nabla_{JU}\nabla f,JV)\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\lambda g_{M}(JU,JV)=0.\end{split}$
Since $(kerF_{*})^{\perp}$ is totally geodesic, $A_{J\eta}=0.$ Therefore, above equation gives
$\begin{split}&\frac{r}{2}g_{N}(\nabla^{N}_{F_{*}JU}F_{*}(\nabla f),F_{*}JV)+\frac{r}{2}g_{N}(\nabla^{N}_{F_{*}JV}F_{*}(\nabla f),F_{*}JU)\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\lambda g_{N}(F_{*}JU,F_{*}JV)=0.\end{split}$
This gives
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla^{N}_{F_{*}JU}F_{*}(\nabla f),F_{*}JV)+g_{N}(\nabla^{N}_{F_{*}JV}F_{*}(\nabla f),F_{*}JU)\Big{)}\\ &+\alpha Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\beta g_{N}(F_{*}JU,F_{*}JV)=0.\end{split}$
Therefore $F_{*}(JkerF_{*})$ is an $\alpha$- Ricci soliton with potential vector field $F_{*}(\nabla f),$ where $\alpha=\frac{1}{r}$ and $\beta=\frac{\lambda}{r}.$
If $\eta\in\Gamma(KerF_{*})^{\perp}$ but $\notin\Gamma(\mu)$, then from (3.18), we have
$\begin{split}&\frac{1}{2}\Big{(}g(T_{U}J\eta,JV)+g(JU,T_{V}J\eta)\Big{)}+\frac{r}{2}g_{M}(\nabla_{JU}\nabla f,JV)+\frac{r}{2}g_{M}(\nabla_{JV}\nabla f,JU)\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\lambda g_{M}(JU,JV)=0.\end{split}$
Using Theorem 2.3 and Theorem 3.5 ([22]) , we get
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla^{N}_{F_{*}JU}F_{*}(\nabla f),F_{*}JV)+g_{N}(\nabla^{N}_{F_{*}JV}F_{*}(\nabla f),F_{*}JU)\Big{)}\\ &+\alpha Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\beta g_{N}(F_{*}JU,F_{*}JV)=0.\end{split}$
This also shows that $F_{*}(JkerF_{*})$ is an $\alpha$- Ricci soliton with potential vector field $F_{*}(\nabla f)$, where $\alpha=\frac{1}{r}$ and $\beta=\frac{\lambda}{r}.$
Corollary 3.3.
Let $F$ be a Clairaut Langrangian Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in(TM)$ and scalar $\lambda$ to a Riemannian manifold $(N,g_{N})$ with $r=e^{f}$. Then $rangeF_{*}$ is an $\alpha$- Ricci soliton.
Proof.
Using Theorem 3.1 and Lemma 56 ([14]), we get the required result. ∎
Theorem 3.2.
Let $F$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to an Einstein manifold $(N,g_{N})$. Then $\nabla f$ is a conformal vector field on $J(kerF_{*}).$
Proof.
Let $F$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold $(M,g_{M},J)$ admitting Ricci soliton with potential vector field $\eta\in\Gamma(TM)$ to an Einstein manifold $(N,g_{N}),$ then we have $L_{\eta}=0$(as we have seen in theorem 3.1). Then from (3.17), we get
$\begin{split}&\frac{r}{2}\Big{(}g_{M}(\nabla_{JU}\nabla f,JV)+g_{M}(\nabla_{JV}\nabla f,JU)\Big{)}\\ &+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)+\lambda g_{M}(JU,JV)=0.\end{split}$ (3.19)
Since $rangeF_{*}$ is Einstein, $Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)=\lambda g_{N}(F_{*}JU,F_{*}JV).$
Then from (2.1) and (3.19), we obtain
$\frac{r}{2}\Big{(}g_{M}(\nabla_{JU}\nabla f,JV)+g_{M}(\nabla_{JV}\nabla f,JU)\Big{)}+2\lambda g_{M}(JU,JV)=0.$
$\frac{1}{2}(L_{\nabla f}g_{M})(JU,JV)+\mu^{\prime}g_{M}(JU,JV)=0,$
where $\mu^{\prime}=\frac{2\lambda}{r}$ is constant on $M.$ This shows that $\nabla f$ is a conformal vector field on $J(kerF_{*}).$
Theorem 3.3.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langrangian Riemannian map from a Kähler manifold admitting Ricci soliton with potential vector fields $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then $s^{(rangeF_{*})}=-rdiv(\nabla f)-\lambda r.$
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with potential vector field $\eta\in\Gamma(TM),$ we have
$\frac{1}{2}(L_{\eta}g_{M})(U,V)+Ric(U,V)+\lambda g_{M}(U,V)=0~{}~{}\forall~{}~{}U,V\in\Gamma(KerF_{*}).\$
If $\eta$ is either horizontal or vertical(as we have seen in theorem 6), we get
$Ric(U,V)+\lambda g_{M}(U,V)=0.$
Using (3.12) in the above equation, we get
$\begin{split}&rg_{M}(\nabla_{JU}\nabla f,JV)+Ric^{(rangeF_{*})}(F_{*}JU,F_{*}JV)\\ &+\lambda g_{M}(U,V)=0.\end{split}$
Taking trace of above equation, we get
$rdiv(\nabla f)+s^{(rangeF_{*})}+\lambda r=0,$
which implies
$s^{(rangeF_{*})}=-rdiv(\nabla f)-\lambda r.$ (3.20)
This completes the proof. ∎
Theorem 3.4.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langragian Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then $kerF_{*}$ is an Einstein.
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$, we have
$\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric(X,Y)+\lambda g_{M}(X,Y)=0.$
Using (3.14) in above equation, we get
$\begin{split}&\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric^{(KerF_{*})}(BX,BY)-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+\lambda g_{M}(X,Y)=0,\end{split}$
which can be rewritten as
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}{\eta},Y)+g_{M}(\nabla_{Y}{\eta},X)\Big{)}-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)\\ &+Ric^{(KerF_{*})}(BX,BY)+\lambda g_{M}(JX,JY)=0.\end{split}$
Since $\eta$ is vertical, above equation gives
$Ric^{(KerF_{*})}(BX,BY)+\nu^{\prime}g_{M}(BX,BY)=0,$ (3.21)
where $\nu^{\prime}=-r|\nabla f|^{2}-div(\nabla f)+\lambda$ is a function on $M.$ This implies that $(kerF_{*})$ is an Einstein.
Corollary 3.4.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut Langragian Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold. Then scalar curvature $s^{kerF_{*}}=-r\nu^{\prime}.$
Proof.
Taking trace of (3.21), we get the result . ∎
Theorem 3.5.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is integrable and $f$ is constant on $\Gamma(\mu).$ Then $kerF_{*}$ is an Einstein, provided $F_{*}(\mu)$ is an Einstein.
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$. Then, we have
$\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric(X,Y)+\lambda g_{M}(X,Y)=0$
which can be rewritten as
$\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric(X,Y)+\lambda g_{M}(X,Y)=0.$
Using Corollary 3.2 in above equation, we get
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-(r|\nabla f|^{2}+div(\nabla f)-\lambda)g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-rCX(f)CY(f)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
Since $(kerF_{*})^{\perp}$ is integrable and $f$ is constant on $\Gamma(\mu),$ from above equation, we get
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)+\lambda g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)+\lambda g_{N}(F_{*}CX,F_{*}CY)=0.\end{split}$
Since $F_{*}(\mu)$ is an Einstein, then above equation reduces to
$\begin{split}Ric^{(kerF_{*})}(BX,BY)+\lambda g_{M}(BX,BY)=0.\end{split}$
Therefore from above equation, we can say that $F_{*}(\mu)$ is an Einstein. ∎
Theorem 3.6.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Kähler manifold admitting Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is integrable and $rangeF_{*}$ is Ricci flat. Then
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)=\Big{(}r|\nabla f|^{2}+div(\nabla f)-\lambda\Big{)}g_{M}(BX,BY)\\ &+rg_{M}(\nabla_{CX}\nabla f,CY)+rCX(f)CY(f)-\lambda g_{M}(CX,CY)=0.\end{split}$ (3.22)
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\eta\in\Gamma(TM),$ we get
$\dfrac{1}{2}(L_{\eta}g_{M})(X,Y)+Ric(X,Y)+\lambda g_{M}(X,Y)=0.$
Making use of Corollary 3.2, we obtain
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-\Big{(}r|\nabla f|^{2}+div(\nabla f)\Big{)}g_{M}(BX,BY)+Ric^{(rangeF_{*})}(F_{*}CX,F_{*}CY)\\ &-rCX(f)CY(f)+\lambda g_{M}(X,Y)=0.\end{split}$
Since $rangeF_{*}$ is Ricci flat, the above equation reduces to
$\begin{split}&\dfrac{1}{2}\Big{(}g_{M}(\nabla_{X}\eta,Y)+g_{M}(\nabla_{Y}\eta,X)\Big{)}+Ric^{(kerF_{*})}(BX,BY)-rg_{M}(\nabla_{CX}\nabla f,CY)\\ &-\Big{(}r|\nabla f|^{2}+div(\nabla f)-\lambda\Big{)}g_{M}(BX,BY)-rCX(f)CY(f)+\lambda g_{M}(CX,CY)=0.\end{split}$
Since $\eta$ is vertical and $(kerF_{*})^{\perp}$ is integrable, we get
$\begin{split}&Ric^{(kerF_{*})}(BX,BY)=\Big{(}r|\nabla f|^{2}+div(\nabla f)-\lambda\Big{)}g_{M}(BX,BY)\\ &+rg_{M}(\nabla_{CX}\nabla f,CY)+rCX(f)CY(f)-\lambda g_{M}(CX,CY).\end{split}$
This completes the proof. ∎
Theorem 3.7.
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a Clairaut anti-invariant Riemannian map from a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\eta\in\Gamma(TM)$ and scalar $\lambda$ to a Riemannian manifold such that $(kerF_{*})^{\perp}$ is integrable. Then
$Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)=\frac{1}{2\alpha^{\prime}}(L_{F_{*}(\nabla f)}g_{N})(F_{*}JU,F_{*}CX)$
where $\alpha^{\prime}=\frac{-1}{r}.$
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\eta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\eta}g_{M})(U,X)+Ric(U,X)+\lambda g_{M}(U,X)=0.$
Making use of (3.2), we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{U}\eta,X)+g_{M}(\nabla_{X}\eta,U)\Big{)}-(r+1)g_{M}(\nabla_{BX}\nabla f,JU)\\ &-rg_{M}(\nabla_{JU}\nabla f,CX)+div\Big{(}A(JU,CX)\Big{)}+Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)\\ &+\sum_{i=1}^{r+s}g_{M}\Big{(}(\nabla_{X_{i}}A)(X_{i},JU),BX)\Big{)}=0.\end{split}$
Since $\eta$ is horizontal and $(kerF_{*})^{\perp}$ is integrable, we get
$\begin{split}-rg_{M}(\nabla_{JU}\nabla f,CX)+Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)=0\end{split}$
which can be rewritten as
$\begin{split}&\frac{1}{2}\Big{(}g_{N}(\nabla_{F_{*}JU}F_{*}(\nabla f),F_{*}CX)+g_{N}(\nabla_{F_{*}{CX}}F_{*}(\nabla f),F_{*}JU)\Big{)}\\ &-\frac{1}{r}Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)=0.\end{split}$
This gives
$\frac{1}{2}(L_{F_{*}(\nabla f)}g_{N})(F_{*}JU,F_{*}CX)+\alpha^{\prime}Ric^{rangeF_{*}}(F_{*}JU,F_{*}CX)=0.$
This proves the required result. ∎
Example 3.1.
Let $(M,g_{M},J)$ be a 6-dimensional Kähler manifold, where $M=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})\in\mathbb{R}^{4}:x_{i}\neq 0,i=1,\ldots,6\},$ with Riemannian metric
$g_{M}=e^{-2x_{3}}dx_{1}^{2}+e^{-2x_{3}}dx_{2}^{2}+dx_{3}^{2}+dx_{4}^{2}+e^{-2x_{6}}dx_{5}^{2}+dx_{6}^{2},$
and $J(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=(-x_{6},-x_{5},-x_{4},-x_{3},-x_{2},-x_{1}).$
Let $(N,g_{N})$ be a 6-dimensional Riemannian manifold, where $N=\{(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6})\in\operatorname{\mathbb{R}}^{6}:y_{j}\in\operatorname{\mathbb{R}}\forall j=1,\ldots,6\}$ with Riemannian metric
$g_{N}=dy_{1}^{2}+dy_{2}^{2}+dy_{3}^{2}+dy_{4}^{2}+dy_{5}^{2}+dy_{6}^{2}.$
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a map defined as
$F(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=(x_{3},0,cosx_{4},sinx_{4},x_{5},x_{6}).$
Then by direct computation, we get
$kerF_{*}=span\Big{\{}U_{1}=e_{1},U_{2}=e_{2}\Big{\}},$
and
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=e_{3},X_{2}=e_{4},X_{3}=e_{5},X_{4}=e_{6}\}.$
Here,
$\{e_{1}=e^{x_{3}}\frac{\partial}{\partial x_{1}},e_{2}=e^{x_{3}}\frac{\partial}{\partial x_{2}},e_{3}=\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}},e_{5}=e^{x_{6}}\frac{\partial}{\partial x_{5}},e_{6}=\frac{\partial}{\partial x_{6}}\}$
is the orthonormal basis of $T_{p}M$, for any $p\in M.$
Then, we obtain
$F_{*}(X_{1})=e_{1}^{\prime},$ $F_{*}(X_{2})=-sinx_{4}e_{3}^{\prime}+cosx_{4}e_{4}^{\prime},F_{*}(X_{3})=e_{5}^{\prime},F_{*}(X_{4})=e_{6}^{\prime},$
where $\{e_{1}^{\prime}=\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}},e_{5}^{\prime}=\frac{\partial}{\partial y_{5}},e_{6}^{\prime}=\frac{\partial}{\partial y_{6}}\}$ represents the standard orthonormal basis of $T_{F(p)}N.$
Also, it can be easily verified that for every $i=1,2,3,4$
$g_{M}(X_{i},X_{i})=g_{N}(F_{*}{X_{i}},F_{*}{X_{i}}),$
and $J(U_{1})=-X_{4},J(U_{2})=-X_{3},J(X_{1})=-X_{2},J(X_{2})=-X_{1},J(X_{3})=-U_{2},J(X_{4})=-U_{1}.$ Thus $F$ is an anti-invariant Riemannian map.
Now in order to show that $F$ is a Clairaut anti-invariant Riemannian map, we need to find a smooth function $f$ on $M$ such that
[…]
For the Riemannian metric $g_{M},$ the Christoffel symbols are obtained as
$\Gamma^{1}_{13}=-1=\Gamma^{1}_{31},\Gamma^{3}_{11}=e^{-2x_{3}},\Gamma^{2}_{23}=-1=\Gamma^{2}_{32},\Gamma^{3}_{22}=e^{-2x_{3}},\Gamma^{5}_{56}=-1=\Gamma^{5}_{65},\Gamma^{6}_{55}=e^{-2x_{6}}$ and rest of them are zero.
Additionally, the covariant derivatives are calculated as
$\nabla_{U_{1}}U_{1}=X_{1},\nabla_{U_{1}}X_{1}=-U_{1},\nabla_{U_{2}}U_{2}=X_{1},\nabla_{U_{2}}X_{1}=-U_{2},\nabla_{X_{3}}X_{3}=X_{4},\nabla_{X_{3}}X_{4}=-X_{3},$ rest of them are zero.
Ricci curvatures are obtained as $Ric(X_{2},X_{2})=-2,Ric(X_{3},X_{3})=-2,Ric(X_{4},X_{4})=-2,$ remaining Ricci tensors are zero.
As we know,
$TM=kerF_{*}\oplus(kerF_{*})^{\perp},$
and for i = $1,\ldots 6$, we have
$Z_{i}=a_{i}U_{1}+b_{i}U_{2}+c_{i}X_{1}+d_{i}X_{2}+l_{i}X_{3}+h_{i}X_{4}.$
Also, we have
$(L_{Z_{1}}g_{M})(Z_{2},Z_{3})=g_{M}(\nabla_{Z_{2}}Z_{1},Z_{3})+g_{M}(\nabla_{Z_{3}}Z_{1},Z_{2}).$
By, utilizing the values of covariant derivatives, we obtain
$\begin{split}(L_{Z_{1}}g_{M})(Z_{2},Z_{3})=&\frac{1}{2}(a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+l_{1}l_{2}h_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+l_{1}l_{3}h_{2})\\ &-a_{2}c_{1}a_{3}-b_{3}b_{2}c_{1}-l_{2}l_{3}h_{1},\end{split}$ (3.23)
$Ric(Z_{2},Z_{3})=-2l_{2}l_{3}-2h_{2}h_{3}-2d_{2}d_{3},$ (3.24)
and
$g_{M}(Z_{2},Z_{3})=a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+l_{2}l_{3}+h_{2}h_{3}.$ (3.25)
Let $(M,g_{M},J)$ admits a Ricci soliton, then
$\frac{1}{2}(L_{Z_{1}}g_{M})(Z_{2},Z_{3})+Ric(Z_{2},Z_{3})+\lambda g_{M}(Z_{2},Z_{3})=0.$
Using (3.23), (3.24) and (3.25) in the above equation, we get
$\begin{split}\lambda=\frac{A_{1}-A_{2}}{2(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+l_{2}l_{3}+h_{2}h_{3})},~{}(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+l_{2}l_{3}+h_{2}h_{3})\neq 0,\end{split}$
where
$A_{1}=2(2l_{2}l_{3}+2h_{2}h_{3}+2d_{2}d_{3}+a_{2}a_{3}c_{1}+b_{2}b_{3}c_{1}+l_{2}l_{3}h_{1}),$
$A_{2}=a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+l_{1}l_{2}h_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+l_{1}l_{3}h_{2}.$
[…]
Example 3.2.
Let $(M,g_{M},J)$ be a 4- dimensional Kähler manifold, where $M=\{(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}:x_{i}\neq 0,i=1,...,4\},$ with Riemannian metric
$g_{M}=e^{-2x_{2}}dx_{1}^{2}+dx_{2}^{2}+e^{-2x_{2}}dx_{3}^{2}+dx_{4}^{2}$
and $J(x_{1},x_{2},x_{3},x_{4})=(-x_{2},x_{1},-x_{4},x_{3}).$ Let $(N,g_{N})$ be a 4-dimensional Riemannian manifold, where $N=\{(y_{1},y_{2},y_{3},y_{4})\in\operatorname{\mathbb{R}}^{4}:y_{j}\in\operatorname{\mathbb{R}}\forall j=1,...,4\}$ with Riemannian metric
$g_{N}=dy_{1}^{2}+dy_{2}^{2}+dy_{3}^{2}+dy_{4}^{2}.$
Let $F:(M,g_{M},J)\rightarrow(N,g_{N})$ be a map, defined as
$F(x_{1},x_{2},x_{3},x_{4})=(cosx_{2},sinx_{2},x_{4},0).$
Then by direct computation, we get
$kerF_{*}=span\Big{\{}U_{1}=e_{1},U_{2}=e_{3}\Big{\}}$
and
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=e_{2},X_{2}=e_{4}\}.$
Here $\{e_{1}=e^{x_{2}}\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=e^{x_{2}}\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}}\}$ is an orthonormal basis of $T_{p}M$, for any $p\in M.$
Then
$F_{*}(X_{1})=-sinx_{2}e_{1}^{\prime}+cosx_{2}e_{2}^{\prime},$ $F_{*}(X_{2})=e_{3}^{\prime},$
where $\{e_{1}^{\prime}=\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}}\}$ is the standard orthonormal basis of $T_{F(p)}N.$ We can easily check that
$g_{M}(X_{i},X_{i})=g_{N}(F_{*}{X_{i}},F_{*}{X_{i}})$
$\forall i=1,2$ and $J(U_{1})=X_{1},J(U_{2})=X_{2}.$ Thus $F$ is a Langrangian Riemannian map.
To show that $F$ is a Clairaut Langragian Riemannian map, we need to find a smooth function $f$ on $M$ such that $T_{U}U=-g(U,U)\nabla f$ for $U\in\Gamma(kerF_{*}).$
For Riemannian metric $g_{M},$ Christoffel symbols are given by
$\Gamma^{2}_{11}=e^{-2x_{2}},\Gamma^{1}_{12}=-1=\Gamma^{1}_{21},\Gamma^{3}_{23}=-1=\Gamma^{3}_{32},\Gamma^{2}_{33}=e^{-2x_{2}}$ and rest of Christoffel symbols are zero.
The covariant derivatives are calculated as
$\nabla_{U_{1}}U_{1}=X_{1},\nabla_{U_{1}}X_{1}=-U_{1},\nabla_{U_{2}}U_{2}=X_{1},\nabla_{U_{2}}X_{1}=-U_{2}$ and rest are zero.
Gradient of any smooth function $f$ is given by
$\nabla f=\sum_{i,j=1}^{4}g^{ij}\frac{\partial f}{\partial x_{i}}\frac{\partial}{\partial x_{j}}.$
Therefore, we get $\nabla f=-X_{1}.$
Using (3.1), (3.2), (3.3) and above covariant derivatives, we get
$Ric(X_{1},X_{1})=-2,Ric(X_{2},X_{2})=-2,$ and rest are zero.
Since $TM=kerF_{*}\oplus(kerF_{*})^{\perp},$ for any $Y_{i}\in\Gamma(TM),Y_{i}=a_{i}U_{1}+b_{i}U_{2}+c_{i}X_{1}+d_{i}X_{2},i=1,...,4,$ we have
$(L_{Y_{1}}g_{M})(Y_{2},Y_{3})=g_{M}(\nabla_{Y_{2}}Y_{1},Y_{3})+g_{M}(\nabla_{Y_{3}}Y_{1},Y_{2}).$
Substituting the values of covariant derivatives, we obtain
$\begin{split}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})=\frac{1}{2}(a_{1}a_{2}b_{3}+c_{2}c_{1}b_{3}+a_{3}a_{1}b_{2}+c_{3}c_{1}b_{2})-a_{2}b_{1}a_{3}-c_{2}b_{1}c_{3},\end{split}$ (3.26)
$Ric(Y_{2},Y_{3})=-2b_{2}b_{3}-2d_{2}d_{3},$ (3.27)
and
$g_{M}(Y_{2},Y_{3})=(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}).$ (3.28)
Let $(M,g_{M},J)$ admits a Ricci soliton, then
$\frac{1}{2}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})+Ric(Y_{2},Y_{3})+\lambda g_{M}(Y_{2},Y_{3})=0.$
Using (3.26), (3.27) and (3.28) in the above equation, we get
$\lambda=\frac{A_{1}-A_{2}}{2(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{2})},~{}a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}\neq 0,$ (3.29)
where
$A_{1}=2(a_{2}b_{1}a_{3}+c_{2}b_{1}c_{3}+2b_{2}b_{3}+2d_{2}d_{3}),$
$A_{2}=a_{1}a_{2}b_{3}+c_{1}c_{2}b_{3}+a_{1}a_{3}b_{2}+c_{1}c_{3}b_{2},$
[…]
4 Clairaut Anti-invariant Riemannian maps to Kähler manifolds admitting Ricci Solitons
In this section, we study Clairaut anti-invariant Riemannian maps from Riemannian manifolds to Kähler manifolds admitting Ricci solitons. Throughout this section, $(rangeF_{*})^{\perp}$ is assumed to be totally geodesic and $dim(rangeF_{*})>1.$
Theorem 4.1.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold with $\tilde{s}=e^{g}$ Then Ricci tensors on $(M,g_{M},J)$ are
$Ric(F_{*}X,F_{*}Y)=Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)$ (4.1)
$\begin{split}Ric(F_{*}X,E)=&Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,QE)+\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{}^{*}F_{*}PE}S)_{JF_{*}X}F_{*}X_{j},F_{*}X_{j}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{X_{j}}}S)_{JF_{*}X}PE,F_{*}X_{j}\Big{)}+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{QE}JF_{*}X)\\ &-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PE,e_{k})JF_{*}X,e_{k}),\end{split}$ (4.2)
$\begin{split}Ric(D,E)=&Ric^{rangeF_{*}}(PD,PE)-\sum_{k=1}^{r_{1}}g_{M}(PD,PE)\Big{(}||\nabla^{M}g||^{2}+H^{g}(e_{k},e_{k})\Big{)}\\ &+\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PD}}S)_{QE}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QE}PD,F_{*}{X_{j}}\Big{)}\\ &-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PD,e_{k})QE,e_{k})+\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{{}^{*}_{F_{*}PE}}S)_{QD}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QD}PE,F_{*}{X_{j}}\Big{)}-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PE,e_{k})QD,e_{k})\\ &+Ric^{(rangeF_{*})^{\perp}}(QD,QE)-(m-r)\Big{(}QD(g)QE(g)+g_{M}(\nabla^{M}_{QD}\nabla^{M}g,QE)\Big{)}\end{split}$ (4.3)
for $X,Y\in\Gamma(kerF_{*})^{\perp},F_{*}X,F_{*}Y\in\Gamma(rangeF_{*})$ and $D,E\in\Gamma(rangeF_{*})^{\perp}$, where $\{F_{*}{X_{j}}\}_{r+1\leq j\leq m}$ and $\{e_{k}\}_{1\leq k\leq n_{1}}$ are orthonormal bases of $rangeF_{*}$ and $(rangeF_{*})^{\perp}$, respectively, where ${}^{*}{F_{*}}$ is the adjoint map of $F_{*}$ and $Ric^{(rangeF_{*})},$ $Ric^{(rangeF_{*})^{\perp}}$ are Ricci tensors on $(rangeF_{*}),$ $(rangeF_{*})^{\perp},$ respectively.
Proof.
Making use of proposition 5.1 [7], we get (4.1) and (4.2). For the proof of (4.3), we proceed as follows.
Since $M$ is a Kähler manifold, we have
$Ric(D,E)=Ric(JD,JE)~{}~{}\forall D,E\in\Gamma(TM).$ (4.4)
Using (2.8), we get
$Ric(JD,JE)=Ric(PD,PE)+Ric(QD,PE)+Ric(PD,QE)+Ric(QD,QE).$ (4.5)
Since
$Ric(PD,PE)=\sum_{j=r+1}^{m}g_{M}(R(F_{*}{X_{j}},PD)PE,F_{*}{X_{j}})+\sum_{k=1}^{n_{1}}g_{M}(R(e_{k},PD)PE,e_{k}).$
Using Theorem 3.1([24]) in above equation, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{n_{1}}g_{M}(S_{\nabla^{F\perp}_{e_{k}}e_{k}}PD,PE)+\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}S_{e_{k}}PD,PE)\\ &-\sum_{k=1}^{n_{1}}g_{M}(S_{e_{k}}PD,S_{e_{k}}PE)-\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}PD,S_{e_{k}}PE).\end{split}$
Using Theorem 3.1([7]) in above equation, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)+\sum_{k=1}^{n_{1}}g_{M}(PD,PE)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{e_{k}}e_{k})\\ &-\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}(e_{k}(g)PD),PE)+\sum_{k=1}^{n_{1}}g_{M}(\nabla^{M}_{e_{k}}PD,e_{k}(g)PE)\\ &-\sum_{k=1}^{n_{1}}(e_{k}(g))^{2}g_{M}(PD,PE).\end{split}$
After simplification, we get
$\begin{split}Ric(PD,PE)&=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{n_{1}}g_{M}(PD,PE)\Big{(}||\nabla^{M}g||^{2}+H^{g}(e_{k},e_{k})\Big{)}.\end{split}$ (4.6)
By similar computation, we get
$\begin{split}Ric(PD,QE)=&\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{*F_{*}PD}S)_{QE}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QE}PD,F_{*}{X_{j}}\Big{)}\\ &-\sum_{k=1}^{n_{1}}g_{M}(R^{F\perp}(PD,e_{k})QE,e_{k}),\end{split}$ (4.7)
$\begin{split}Ric(PE,QD)=&\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{*F_{*}PE}S)_{QD}F_{*}{X_{j}},F_{*}{X_{j}}\Big{)}\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{QD}PE,F_{*}{X_{j}}\Big{)}-\sum_{k=1}^{n_{1}}g_{M}(R^{F\perp}(PE,e_{k})QD,e_{k})\end{split}$ (4.8)
and
$\begin{split}Ric(QD,QE)=Ric^{(rangeF_{*})^{\perp}}(QD,QE)-(m-r)\Big{(}QD(g)QE(g)+g_{N}(\nabla^{M}_{QD}\nabla^{M}g,QE)\Big{)}.\end{split}$ (4.9)
Adding (4.6), (4.7), (4.8) and (4.9), we get (4.5) and with the help of (4.4), we get (4.3). ∎
Corollary 4.1.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langragian Riemannian map from a Riemannian manifold to a Kähler manifold with $\tilde{s}=e^{g}$. Then Ricci curvature is given by
$Ric(F_{*}X,F_{*}Y)=Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)$ (4.10)
$\begin{split}Ric(F_{*}X,E)=&\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{*_{F_{*}PE}}S)_{JF_{*}X}F_{*}X_{j},F_{*}X_{j})\\ &-\sum_{j=r+1}^{m}g_{M}\Big{(}(\tilde{\nabla}_{X_{j}}S)_{JF_{*}X}PE,F_{*}X_{j})\\ &-\sum_{k=1}^{r_{1}}g_{M}(R^{F\perp}(PE,e_{k})JF_{*}X,e_{k}),\end{split}$ (4.11)
$\begin{split}Ric(D,E)=Ric^{(rangeF_{*})}(PD,PE)-\sum_{k=1}^{r_{1}}g_{M}(PD,PE)\Big{(}(||\nabla^{M}g||^{2}+H^{g}(e_{k},e_{k})\Big{)},\end{split}$ (4.12)
where $D,E\in\Gamma({rangeF_{*}})^{\perp}.$
Proof.
In case of Langrangian Riemannian map, $\nu=0,$ this implies $QD=QE=0.$ Applying this in (4.1), (4.2) and (4.3), we get the required result. ∎
Theorem 4.2.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with vertical potential vector field $\vartheta$ and scalar $\lambda$ such that $(rangeF_{*})^{\perp}$ is an Einstein. Then $\vartheta$ is a conformal vector field on $rangeF_{*}$.
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\vartheta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric(F_{*}X,F_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.$
Making use of (4.1) in above equation, we obtain
$\begin{split}&\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0,\end{split}$
which reduces to
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{F_{*}X}\vartheta,F_{*}Y)+g_{M}(\nabla_{F_{*}Y}\vartheta,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$
Since $F$ is a totally geodesic Clairaut anti-invariant Riemannian map, second fundamental form of $F$ vanishes. Then by Theorem 3.2 ([7]), g is constant on $M.$ Therefore, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{N}_{F_{*}X}\vartheta,F_{*}Y)+g_{M}(\nabla^{N}_{F_{*}Y}\vartheta,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$ (4.13)
Since $\vartheta$ is vertical, i.e., $\vartheta=F_{*}Z,$ then from above equation, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{M}_{F_{*}X}F_{*}Z,F_{*}Y)+g_{M}(\nabla^{M}_{F_{*}Y}F_{*}Z,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+\lambda g_{M}(JF_{*}X,JF_{*}Y)=0.\end{split}$ (4.14)
Since $(rangeF_{*})^{\perp}$ is an Einstein,
$Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)=\lambda g_{M}(JF_{*}X,JF_{*}Y)=\lambda g_{M}(F_{*}X,F_{*}Y),$
using above equation in (4.14), we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{M}_{F_{*}X}F_{*}Z,F_{*}Y)+g_{M}(\nabla^{M}_{F_{*}Y}F_{*}Z,F_{*}X)\Big{)}+2\lambda g_{M}(F_{*}X,F_{*}Y)=0,\end{split}$
this gives
$\frac{1}{2}(L_{F_{*}Z}g_{M})(F_{*}X,F_{*}Y)+2\lambda g_{M}(F_{*}X,F_{*}Y)=0.$
This implies that $F_{*}Z$ is conformal vector field on $(rangeF_{*}).$
Theorem 4.3.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a totally geodesic Clairaut anti-invariant Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta$ and scalar $\lambda.$ Then $J(rangeF_{*})$ is an Einstein.
Proof.
Since $\vartheta$ is horizontal vector field, i.e., $\vartheta=V,$ then from (4.13), we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla^{M}_{F_{*}X}V,F_{*}Y)+g_{M}(\nabla^{M}_{F_{*}Y}V,F_{*}X)\Big{)}+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$
Using (2.4) and Theorem 3.1([7]), we get
$V(g)g_{M}(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+\lambda g_{M}(JF_{*}X,JF_{*}Y)=0.$
This implies
$Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+\rho g_{M}(JF_{*}X,JF_{*}Y)=0,$
where $\rho=V(g)+\lambda.$ This proves that $J(rangeF_{*})$ is an Einstein. ∎
Corollary 4.2.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langragian Riemannian map from a Riemannian manfold $(N,g_{N})$ to a Kähler manifold $(M,g_{M},J),$ admitting Ricci soliton with horizontal potential vector field $\vartheta$ and scalar $\lambda$ then, $(rangeF_{*})^{\perp}$ is an Einstein.
Proof.
Since $F$ is a Langragian Riemannian map, $J(rangeF_{*})=(rangeF_{*})^{\perp}.$ Using Theorem (4.3), we see that $(rangeF_{*})^{\perp}$ is an Einstein. ∎
Theorem 4.4.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langragian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TM)$ and scalar $\lambda.$ Then scalar curvature $s^{(rangeF_{*})^{\perp}}$ of $(rangeF_{*})^{\perp}$ is given by
$s^{(rangeF_{*})^{\perp}}=-\rho(m-r),$ (4.15)
where $\rho=V(g)+\lambda.$
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric(F_{*}X,F_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.$
Using (4.1) in above equation, we get
$\begin{split}&\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)+Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)\\ &+(m-r)g_{M}(\nabla^{M}g,\nabla^{F\perp}_{JF_{*}X}JF_{*}Y)+\lambda g_{M}(F_{*}X,F_{*}Y)=0.\end{split}$
In Theorem 4.3, we have seen that $\frac{1}{2}(L_{\vartheta}g_{M})(F_{*}X,F_{*}Y)$ = $V(g)g_{M}(F_{*}X,F_{*}Y),$ using Theorem 5.2 [7], we get
$\begin{split}&Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)+\rho g_{M}(F_{*}X,F_{*}Y)=0,\end{split}$
which gives
$\begin{split}Ric^{(rangeF_{*})^{\perp}}(JF_{*}X,JF_{*}Y)=-\rho g_{M}(JF_{*}X,JF_{*}Y).\end{split}$
Taking trace of above equation, we get
$\begin{split}s^{(rangeF_{*})^{\perp}}=-\rho n_{1}\end{split}$
This completes the proof. ∎
Theorem 4.5.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with vertical potential vector field $\vartheta$ and scalar $\lambda.$ Then $rangeF_{*}$ is an Einstein.
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with vertical potential vector field $\vartheta\in\Gamma(TM)$, we have
$\frac{1}{2}(L_{\vartheta}g_{M})(D,E)+Ric(D,E)+\lambda g_{M}(D,E)=0.$
Using (4.12) in above equation, we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{D}\vartheta,E)+g_{M}(\nabla_{E}\vartheta,D)+Ric^{(rangeF_{*})}(PD,PE)\\ &+\sum_{k=1}^{r_{1}}g_{M}(PD,PE)\Big{(}-H^{g}(e_{k},e_{k})-||\nabla^{M}g||^{2}+\lambda\Big{)}=0.\end{split}$ (4.16)
Since $\vartheta$ is vertical, i.e., $\vartheta=F_{*}X,$ we get
$\begin{split}&\frac{1}{2}\Big{(}g_{M}(\nabla_{D}F_{*}X,E)+g_{M}(\nabla_{D}F_{*}X,E)+Ric^{(rangeF_{*})}(PD,PE)\\ &+\sum_{k=1}^{r_{1}}g_{M}(PD,PE)(\lambda-H^{g}(e_{k},e_{k})-||\nabla^{M}g||^{2})=0.\end{split}$
Using metric compatibility, we obtain
$Ric^{(rangeF_{*})}+\psi g_{M}(PD,PE)=0,$ (4.17)
where $\psi=\sum_{k=1}^{r_{1}}(\lambda-H^{g}(e_{k},e_{k})-||\nabla^{M}g||^{2})$ is a function on $M$. This shows that $rangeF_{*}$ is an Einstein. ∎
Theorem 4.6.
Let $F:(N,g_{N})\rightarrow(M,g_{M},J)$ be a Clairaut Langangian Riemannian map from a Riemannian manifold to a Kähler manifold admitting Ricci soliton with horizontal potential vector field $\vartheta$ and a scalar $\lambda.$ Then the scalar curvature $s^{(rangeF_{*})^{\perp}}$ of $(rangeF_{*})^{\perp}$ is given by $s^{(rangeF_{*})^{\perp}}=-r\psi.$
Proof.
Since $(M,g_{M},J)$ is a Kähler manifold admitting a Ricci soliton with horizontal potential vector field $\vartheta\in\Gamma(TM)$, we have
$\begin{split}\frac{1}{2}(L_{\vartheta}g_{M})(D,E)+Ric(D,E)+\lambda g_{M}(D,E)=0.\end{split}$
Using (4.12) and since $\vartheta$ is horizontal vector field, i.e., $\vartheta=W,$ we have
$\begin{split}\frac{1}{2}\Big{(}g_{M}(\nabla_{D}W,E)+g_{M}(D,\nabla_{E}W)\Big{)}+Ric^{(rangeF_{*})}(PD,PE)+\psi g_{M}(PD,PE)=0.\end{split}$
Taking trace of above equation, we get
$\sum_{k=1}^{r_{1}}g_{M}(\nabla_{e_{k}}e_{k},e_{k})+\sum_{j=r+1}^{m}Ric^{(rangeF_{*})}(F_{*}{X_{j}},F_{*}{X_{j}})+\sum_{j=r+1}^{m}\psi g_{M}(F_{*}{X_{j}},F_{*}{X_{j}})=0$
which gives
$s^{(rangeF_{*})}=-r\psi.$ (4.18)
If $\vartheta$ is vertical vector field, then taking trace of (4.17), we get the same result. This completes the proof. ∎
Example 4.1.
Let $M=\{(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6}):x_{i}\neq 0,i=1,...,6\}$ be a Riemannian manifold with Riemannian metric $g_{M}=dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}+dx_{4}^{2}+dx_{5}^{2}+dx_{6}^{2}$ and $N=R^{6}$ be a Kähler manifold with Riemannian metric and complex structure $g_{N}=e^{-2y_{4}}dy_{1}^{2}+dy_{2}^{2}+e^{-2y_{4}}dy_{3}^{2}+dy_{4}^{2}+dy_{5}^{2}+dy_{6}^{2},$ $J(a,b,c,d,l,h)=(-b,a,-d,c,-h,l)$ respectively. We construct a map $F:(M,g_{M})\rightarrow(N,g_{N},J),$ defined by
$F(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=\big{(}\frac{x_{1}-x_{4}}{\sqrt{2}},0,x_{2},0,0,0\Big{)}$
Then $kerF_{*}=span\{U_{1}=\frac{e_{1}+e_{4}}{\sqrt{2}},U_{2}=e_{3},U_{3}=e_{5},U_{4}=e_{6}\}$ and $(kerF_{*})^{\perp}=span\{X_{1}=\frac{e_{1}-e_{4}}{\sqrt{2}},X_{2}=e_{2}\},$ where $\{e_{1}=\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}},e_{5}=\frac{\partial}{\partial x_{5}},e_{6}=\frac{\partial}{\partial x_{6}}\}$ and $\{e_{1}^{\prime}=e^{y_{4}}\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=e^{y_{4}}\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=\frac{\partial}{\partial y_{4}},e_{5}^{\prime}=\frac{\partial}{\partial y_{5}},e_{6}^{\prime}=\frac{\partial}{\partial y_{6}}\}$ are the bases of $T_{p}M$ and $T_{F(p)N}$ respectively.
Here $F_{*}(X_{1})=e_{1}^{\prime}$ and $F_{*}(X_{2})=e_{3}^{\prime}$ and we can easily compute that
$g_{M}(X,X)=g_{N}(F_{*}X,F_{*}X)~{}~{}\forall~{}X\in(kerF_{*})^{\perp}.$
$rangeF_{*}=span\{F_{*}{X_{1}}=e_{1}^{\prime},F_{*}{X_{2}}=e_{3}^{\prime}\}$ and $(rangeF_{*})^{\perp}=\{e_{2}^{\prime},e_{4}^{\prime},e_{5}^{\prime},e_{6}^{\prime}\}.$
Also we can see that $JF_{*}{X_{1}}=e_{2}^{\prime},JF_{*}{X_{2}}=e_{4}^{\prime}$ means $F$ is an anti-invariant Riemannian map.
To show $F$ is a Clairaut anti-invariant Riemanian map, we have to find $g$ such that
$(\nabla F_{*})(X,X)=-g(X,X)\nabla^{M}g.$
After some computation, we get
$\Gamma^{1}_{14}=-1=\Gamma^{1}_{41},\Gamma^{3}_{34}=-1=\Gamma^{3}_{43},\Gamma^{4}_{11}=e^{-2y_{4}},\Gamma^{4}_{33}=e^{-2y_{4}}$ and rest of Christoffel symbols are zero.
The covariant derivative is given as
$\nabla_{F_{*}{X_{1}}}{F_{*}{X_{1}}}=e_{4}^{\prime},\nabla_{F_{*}{X_{1}}}e_{4}^{\prime}=-e_{1}^{\prime},\nabla_{F_{*}{X_{2}}}{F_{*}{X_{2}}}=e_{4}^{\prime},\nabla_{F_{*}{X_{2}}}e_{4}^{\prime}=-e_{3}^{\prime}$ and rest of covariant derivatives are zero.
Gradient of any smooth function $g$ is given by
$\nabla^{M}g=\sum_{i,j=1}^{6}g^{ij}\frac{\partial g}{\partial y_{i}}\frac{\partial}{\partial y_{j}}.$
Therefore, we get $\nabla^{M}g=-e_{4}^{\prime}.$
Since $TM=rangeF_{*}\oplus(rangeF_{*})^{\perp},$ for any $T_{i}=a_{i}F_{*}X_{1}+b_{i}e_{2}^{\prime}+c_{i}F_{*}X_{2}+d_{i}e_{4}^{\prime}+h_{i}e_{5}^{\prime}+l_{i}e_{6}^{\prime},i=1,...,6.$
Since
$(L_{T_{1}}g_{M})(T_{2},T_{3})=g_{M}(\nabla_{T_{2}}T_{1},T_{3})+g_{M}(\nabla_{T_{3}}T_{1},T_{2}).$
Substituting the values of covariant derivatives, we obtain
$\begin{split}\frac{1}{2}(L_{T_{1}}g_{M})(T_{2},T_{3})=\frac{1}{2}\Big{(}a_{1}a_{2}d_{3}-a_{2}d_{1}a_{3}+c_{1}c_{2}d_{3}-c_{2}c_{3}d_{1}+a_{1}a_{3}d_{2}-a_{3}d_{1}a_{2}+c_{3}c_{1}d_{2}-c_{3}d_{1}c_{2}\Big{)},\end{split}$
$g_{M}(T_{2},T_{3})=(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+h_{2}h_{3}+l_{2}l_{3}),$
$Ric(T_{2},T_{3})=-a_{2}a_{3}-d_{2}d_{3}.$
Let $(M,g_{M},J)$ admits a Ricci soliton, then
$\frac{1}{2}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})+Ric(Y_{2},Y_{3})+\lambda g_{M}(Y_{2},Y_{3})=0$
which implies
$\lambda=\frac{a_{2}a_{3}+d_{2}d_{3}+a_{2}a_{3}d_{1}+c_{2}c_{3}d_{1}-\frac{1}{2}\Big{(}a_{1}a_{2}d_{3}+c_{1}c_{2}d_{3}+a_{1}a_{3}d_{2}+c_{3}c_{1}d_{2}\Big{)}}{(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+h_{2}h_{3}+l_{2}l_{3})},$
where $a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}+h_{2}h_{3}+l_{2}l_{3}\neq 0.$ For the suitable choice of scalar’s viz, $a_{i},b_{i},c_{,}d_{i},h_{i},l_{i}$ the behaviour of Ricci soliton will be expanding, steady, shrinking according to $\lambda>0,$ $\lambda=0,$ $\lambda<0$.
Example 4.2.
Let $M=\{(x_{1},x_{2},x_{3},x_{4}):x_{i}\neq 0,\forall i=1,...,4\}$ be a Riemannian manifold with Riemannian metric $g_{M}=dx_{1}^{2}+dx_{2}^{2}+dx_{3}^{2}+dx_{4}^{2}$ and $N=R^{4}$ be a Kähler manifold with Riemannian metric $g_{N}=e^{-2y_{3}}dy_{1}^{2}+e^{-2y_{3}}dy_{2}^{2}+dy_{3}^{2}+e^{-2y_{3}}dy_{4}$ and complex structure $J(a,b,c,d)=(-d,-c,-b,-a).$ Now we construct a map $F:(M,g_{M})\rightarrow(N,g_{N},J),$ defined by
$F(x_{1},x_{2},x_{3},x_{4})=\Big{(}\frac{x_{1}+x_{2}}{\sqrt{2}},\frac{x_{3}+x_{4}}{\sqrt{2}},0,0\Big{)}.$
After some calculations, we get
$kerF_{*}=span\Big{\{}U_{1}=\frac{e_{1}-e_{2}}{\sqrt{2}},U_{2}=\frac{e_{3}-e_{4}}{\sqrt{2}}\Big{\}}$
and
$(kerF_{*})^{\perp}=span\Big{\{}X_{1}=\frac{e_{1}+e_{2}}{\sqrt{2}},X_{2}=\frac{e_{3}+e_{4}}{\sqrt{2}}\Big{\}},$
where $\{e_{1}=\frac{\partial}{\partial x_{1}},e_{2}=\frac{\partial}{\partial x_{2}},e_{3}=\frac{\partial}{\partial x_{3}},e_{4}=\frac{\partial}{\partial x_{4}}\}$ and $\{e_{1}^{\prime}=e^{y_{3}}\frac{\partial}{\partial y_{1}},e_{2}^{\prime}=e^{y_{3}}\frac{\partial}{\partial y_{2}},e_{3}^{\prime}=\frac{\partial}{\partial y_{3}},e_{4}^{\prime}=e^{y_{3}}\frac{\partial}{\partial y_{4}}\}$ are the bases of $T_{p}M$ and $T_{F(p)N}$ respectively.
Here, $F_{*}(X_{1})=e_{1}^{\prime},F_{*}(X_{2})=e_{2}^{\prime}$ and we can easily compute that
$g_{M}(X_{i},X_{i})=g_{N}(F_{*}{X_{i}},F_{*}{X_{i}})~{}~{}\forall X_{i}\in\Gamma(TM).$
$rangeF_{*}=span\{F_{*}(X_{1})=e_{1}^{\prime},F_{*}(X_{2})=e_{2}^{\prime}\}$ and $(rangeF_{*})^{\perp}=\{e_{3}^{\prime},e_{4}^{\prime}\}.$
Also we can see that $JF_{*}(X_{1})=-e_{4}^{\prime},JF_{*}{X_{2}}=-e_{3}^{\prime},$ means $F$ is a Langrarian Riemannian map.
To show that $F$ is a Clairaut Langragian Riemanian map, we have to find $g$ such that
$(\nabla F_{*})(X,X)=-g(X,X)\nabla^{M}g.$
After some computation, we get that for Riemannian metric $g_{M},$ Christoffel symbols are given by
$\Gamma^{1}_{13}=-1=\Gamma^{1}_{31},\Gamma^{2}_{23}=-1=\Gamma^{2}_{32},\Gamma^{3}_{11}=e^{-2y_{3}},\Gamma^{3}_{22}=e^{-2y_{3}},\Gamma^{3}_{44}=e^{-2y_{3}},\Gamma^{4}_{34}=-1=\Gamma^{4}_{43}$ and rest of Christoffel symbols are zero.
The covariant derivative are calculated as
$\nabla_{F_{*}{X_{1}}}F_{*}{X_{1}}=e_{3}^{\prime},\nabla_{F_{*}{X_{1}}}e_{3}^{\prime}=-e_{1}^{\prime},\nabla_{F_{*}{X_{2}}}F_{*}{X_{2}}=e_{3}^{\prime},\nabla_{F_{*}{X_{2}}}e_{3}^{\prime}=-e_{2}^{\prime},\nabla_{e_{4}^{\prime}}e_{3}^{\prime}=-e_{4}^{\prime},\nabla_{e_{4}^{\prime}}e_{4}^{\prime}=e_{3}^{\prime}$and rest of covariant derivatives are zero.
Gradient of any smooth function $g$ is given by
$\nabla^{M}g=\sum_{i,j=1}^{4}g^{ij}\frac{\partial g}{\partial y_{i}}\frac{\partial}{\partial y_{j}}.$
Therefore, we get $\nabla^{M}g=-e_{3}^{\prime}.$
After some calculations, using (3.1), (3.2), (3.3) and above covariant derivatives, we get
$Ric(F_{*}X_{1},F_{*}X_{1})=-2,Ric(e_{3}^{\prime},e_{3}^{\prime})=-2,Ric(e_{4}^{\prime},e_{4}^{\prime})=-2$ and rest are zero.
Since $TM=rangeF_{*}\oplus(rangeF_{*})^{\perp},$ for any $T_{i}=a_{i}F_{*}X_{1}+b_{i}F_{*}X_{2}+c_{i}e_{3}^{\prime}+d_{i}e_{4}^{\prime},i=1,...,4.$
Since
$(L_{T_{1}}g_{M})(T_{2},T_{3})=g_{M}(\nabla_{T_{2}}T_{1},T_{3})+g_{M}(\nabla_{T_{3}}T_{1},T_{2}).$
Substituting the values of covariant derivatives in the above equation, we obtain
$\begin{split}\frac{1}{2}(L_{T_{1}}g_{M})(T_{2},T_{3})=&\frac{1}{2}\Big{(}a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+d_{1}d_{2}c_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+d_{3}d_{1}c_{2}\Big{)}\\ &-a_{2}c_{1}a_{3}-b_{2}b_{3}c_{1}-d_{2}d_{3}c_{1},\end{split}$
$Ric(T_{2},T_{3})=-2a_{2}a_{3}-2c_{2}c_{3}-2d_{2}d_{3},$
and
$g_{M}(T_{2},T_{3})=(a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}).$
We know that $(M,g_{M},J)$ admits a Ricci soliton, if
$\frac{1}{2}(L_{Y_{1}}g_{M})(Y_{2},Y_{3})+Ric(Y_{2},Y_{3})+\lambda g_{M}(Y_{2},Y_{3})=0$
this implies
$\lambda=\frac{P_{1}-P_{2}}{a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}},a_{2}a_{3}+b_{2}b_{3}+c_{2}c_{3}+d_{2}d_{3}\neq 0.$
where
$P_{1}=a_{2}a_{3}(c_{1}+2)+d_{2}d_{3}(c_{1}+2)+b_{2}b_{3}c_{1}+2c_{2}c_{3},$
$P_{2}=\frac{1}{2}\Big{(}a_{1}a_{2}c_{3}+b_{1}b_{2}c_{3}+d_{1}d_{2}c_{3}+a_{1}a_{3}c_{2}+b_{1}b_{3}c_{2}+d_{3}d_{1}c_{2}\Big{)},$
for the suitable choice of scalar’s viz, $a_{i},b_{i},c_{,}d_{i}\in\operatorname{\mathbb{R}},$ the behaviour of Ricci soliton will be expanding, steady, shrinking according to $\lambda>0,$ $\lambda=0,$ $\lambda<0$.
5 Acknowledgement
First author is thankful to the Department of Science and Technology(DST) Government of India for providing financial assistance in terms of FIST project(TPN-69301) vide the letter with Ref No.:(SR/FST/MS-1/2021/104). The second author is grateful to the financial support provided by CSIR (Council of science and industrial research) Delhi, India. File no.[09/1051(12062)/2021-EMR-I].